OpenStudy (anonymous):

the equation of line EF is y=1/2+6. write an equation of a line parallel to line EF in slope-intercept form that contains point (0,2)

OpenStudy (caozeyuan):

I think x is missing

OpenStudy (caozeyuan):

because your eqn is y=13/2

OpenStudy (caozeyuan):

actually this makes life so much easier

OpenStudy (caozeyuan):

since you have a line with 0 slope and it pass (0,2), IT HAS TO BE Y=2

OpenStudy (anonymous):

The options are y=-2x-2 Y=-1/2x+2 Y=1/2x-2 Y=-2x+2

OpenStudy (caozeyuan):

ok so what is your slope of the original line?

OpenStudy (anonymous):

2?

OpenStudy (caozeyuan):

no, it is 1/2

OpenStudy (caozeyuan):

your line is in the form of y=kx+b where k is the slope

OpenStudy (anonymous):

Ohhb yeah your right so C?

OpenStudy (caozeyuan):

No

OpenStudy (caozeyuan):

Wait, do you made a typo? because none of the options are correct

OpenStudy (anonymous):

Ohhhh its (0,-2) on the question **** so sorry

OpenStudy (anonymous):

Which makes C the right answer right?

OpenStudy (caozeyuan):

GREAT!

OpenStudy (anonymous):

Is that correct?

OpenStudy (caozeyuan):

YES, definitely